# Exploring Lee-Yang and Fisher Zeros in the 2D Ising Model through Multi-Point Padé Approximants

### v2

@@ V.1 Extracting $\nu$ and $\beta_{c}$
Figure 9: (Top) The scaling in $1/L$ of $\mbox{Im}(\beta_{0})$, i.e. the imaginary part of the Fisher zero, detected as the closest singularity of the cumulant to the real axis. The correct critical exponent $\nu=1$ is reproduced with fairly good accuracy. (Bottom) Once $\nu$ has been extracted from the data, one can fit the value of the critical inverse temperature $\beta_{c}$, which is reconstructed to per mille accuracy. Note that $\kappa_{1,2}$ are some parameters which encode finite volume effects.
@@
    with $A$ being the logarithm of the proportionality constant in Eq. (10) and $B$ is the exponent of $\frac{1}{L}$ we want to extract and compare its value with $\nu$. As can be seen, the value of the relevant critical exponent $\nu$ is obtained with decent accuracy with a value of $1.014(60)$, its exact value being $\nu=1$ [29]. Shown in the top panel of Fig. 9 is a pictorial description of the fit described in Eq. (12). We plot $\mbox{Im}(\beta_{0})$ as a function of $1/L^{1/B}$, which on account of $B\sim 1$ looks like a straight line.
  • Using Eqs. (10) and (11) and the fact the exponent $\nu=1$, it is not hard to see that one can obtain a linear relation for Re[$\beta_{0}$] as a function of $1/L$ and define a fit function to extract $\beta_{c}$ as follows
    Fit $\textrm{{Fit II}}:\textrm{Re}[\beta_{0}]=A+B*\frac{1}{L}$ (13)
    Here $\nu=1$ simplifies Eq. (11) with $\textrm{Re}[\beta_{0}]\to\beta_{c}$ in the limit $L\to\infty$. Hence, in order to determine $\beta_{c}$, we will fit the real part of the closest Fisher zeros to the real $\beta$ axis as a function of $1/L$ and extract the intercept $A$. This intercept is shown as a green star in the bottom panel of Fig. 9. Also for this our estimate seems fairly accurate at $\beta_{c}=0.4405(5)$, when compared with the exact result of $\beta_{c}\sim 0.4407$ [29]. We additionally show in the same figure, that after identifying the exponent, one can also find the intercept of the line $\mbox{Im}(\beta_{0})$ vs $1/L$ and show that it goes to zero within errors as seen with the red star on the figure.
@@ V.2 Extracting $\beta\delta$
Remembering the value of $\nu=1$ for the 2D Ising model, we can thus use the fit result for the parameter $B$ to estimate $\beta\delta$. As can be seen from Table 1, the fit parameter $B=-1.879(16)$ has been determined, which gives to very good precision the estimate for $\beta\delta$ whose exact value for the 2D Ising model is $1.875$. Further, without using the hyperscaling relation given in Eq. (16), the fit parameter $B$ should be compared against $\beta/\nu-d$, to obtain $\beta=0.121(16)$ which compared against its exact value of $0.125$ for the 2D Ising model, gives a decent estimate for the exponent. Finally, we show the results of the fit of $H_{0}$ we obtained for each lattice size, plotted against $L^{\beta/\nu-d}$ in Fig. 10.
Figure 10: Finite size scaling of $\mbox{Im}(H_{0})$. To guide the eye, we plot data versus $L^{1/8-2}$, where the correct critical exponents $\beta$ and $\nu$ are taken. The value obtained from fits is $\beta/\nu-d=-\beta\delta=-1.879(16)$, as shown in Table 1, which also gives $\beta=0.121(16)$.
In principle one should be able to follow these steps to estimate the critical region for QCD using Taylor expansions from lattice QCD. However, being a much more complicated theory, both numerically and conceptually, we may have to wait for some time to be able to do that.
@@
II $\mathbf{0.4405(5)}$ $-0.2098(77)$ $\beta_{c}\sim 0.4407$ $0.1$
III $1.294(44)$ $\mathbf{1.879(16)}$ $d-\beta/\nu=\beta\delta=1.875$ $3.9$
Table 1: Results for the fits shown in Eqs. (12,13,15), shown for each row respectively. For Fit I, $B$ has to be compared with the exact value stated, whereas for Fit II, the intercept gives $\beta_{c}$, hence the exact value has to be compared with $A$. For the last fit, Fit III, the fit parameter $B$ has to be compared to the exact value of the critical exponent product, namely, $\beta\delta$.
@@ VI Conclusions and Outlook
As a first step we simulated the 2D Ising model using a cluster spin flip algorithm in two ways. For the LY zero analysis, we simulated the model on varying lattice sizes at a set of values of the external magnetic field. These simulations were performed at $T_{c}$, and the goal was to approximate the average magnetization as a rational function of the external magnetic field and study the structure of zeros and poles that arise. On the one hand we were able to verify numerically, many properties of the LY zeros including the famous circle theorem for the Ising model, observing that the only genuine poles (un-cancelled and stable) of the magnetization lie on the purely imaginary $H$ axis. It was further observed that the number of genuine poles poles increases with volume and for simulations at $T_{c}$, comes closer to the real $H$ axis. In order to verify that these were indeed physical effects, volume scaling of the zeros, using the prescription in [11, 17] were performed leading to very precise value of the combination of critical exponents $\beta\delta=1.879(16)$. The fact that LY zeros can be studied at $T_{c}$ is in our approach fully self-consistent. In fact, Fisher zeros were also studied by approximating the specific heat with a multi-point Padé function and studying its poles in the complex $\beta$ plane. Finite size scaling of these zeros following the prescription of [17] was done to obtain again, precise values of the critical exponent $\nu=1.014(60)$ and of the critical inverse temperature $\beta_{c}=0.4405(5)$. These results give us some confidence in our pursuits of studying the QCD phase diagram using lattice QCD simulations combined with multi-point Padé method. The main caveat being that in the case of the Ising model it was not very computationally expensive to reach statistics of the order of $\sim 650$K configurations for each lattice, for each value of $H$ and $\beta$. These kind of statistics are not currently realistic for lattice QCD simulations.
Another direction to pursue would be to study the universal location of the Lee-Yang edge singularities as done using the FRG approach in [57, 58] or using a suitable parametrization of lattice data as done in [59] for the $O(N)$ models. In order to do this using our approach, we would need accurate values of the closest LY zero to the real $H$ axis, but at temperatures $T>T_{c}$. We would like to close by inviting the reader to use the multi-point Padé method for extracting Lee-Yang and Fisher zeros of their choice of models.
@@ Acknowledgements
S.S. would like to thank Christian Schmidt, Frithjof Karsch and Vladimir Skokov for interesting discussions on the critical phenomena and Lee-Yang zeros. This work was supported by the (i) European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie Grant Agreement No. H2020-MSCAITN-2018-813942 (EuroPLEx), (ii) Part of this work was done with S.S. being supported by PUNCH4NFDI consortium supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 460248186 (PUNCH4NFDI), (iii) I.N.F.N. under the re- search project i.s. QCDLAT. A good fraction of this research used computing resources made available through the University of Parma on the UNIPR HPC facility.

### v3

@@ V.1 Extracting $\nu$ and $\beta_{c}$
Refer to captionRefer to caption
[…]
@@
    with $A$ being the logarithm of the proportionality constant in Eq. (10) and $B$ is the exponent of $\frac{1}{L}$ that we want to extract and compare its value with $\nu$. As can be seen in Table. 1, the value of the relevant critical exponent $\nu$ is obtained with decent accuracy with a value of $1.014(60)$, its exact value being $\nu=1$ [35]. Shown in the top panel of Fig. 9 is a pictorial description of the fit described in Eq. (12). We plot $\mbox{Im}(\beta_{0})$ as a function of $1/L$. On account of $B\sim 1$, the dash-dotted line (which is the result of the fit) can hardly be distinguished from a straight line.
  • Using Eqs. (10) and (11) and the fact that the exponent $\nu=1$, it is not hard to see that one can obtain a linear relation for Re[$\beta_{0}$] as a function of $1/L$ and define a fit function to extract $\beta_{c}$ as follows
    Fit $\textrm{{Fit IIa}}:\textrm{Re}[\beta_{0}]=A+B*\frac{1}{L}$ (13)
    Fit $\textrm{{Fit IIb}}:\textrm{Im}[\beta_{0}]=C+D*\frac{1}{L}$ (14)
    Here $\nu=1$ simplifies Eq. (11) with $\textrm{Re}[\beta_{0}]\to\beta_{c}$ in the limit $L\to\infty$. Hence, in order to determine $\beta_{c}$, we will fit the real part of the closest Fisher zeros to the real $\beta$ axis as a function of $1/L$ and extract the intercept $A$. This intercept is shown as a green star in the bottom panel of Fig. 9. Also for this our estimate seems fairly accurate at $\beta_{c}=0.4404(19)$, when compared with the exact result of $\beta_{c}\sim 0.4407$ [35]. We additionally show in the same figure, that after identifying the exponent, one can also find the intercept of the line $\mbox{Im}(\beta_{0})$ vs $1/L$ and show that it goes to zero within errors as seen with the red star on the figure.
@@ V.2 Extracting $\beta\delta$
Remembering the value of $\nu=1$ for the 2D Ising model, we can thus use the fit result for the parameter $B$ to estimate $\beta\delta$. As can be seen from Table 1, the fit parameter $B=-1.881(93)$ has been determined, which gives to decent precision the estimate for $\beta\delta$ whose exact value for the 2D Ising model is $1.875$. Further, without using the hyperscaling relation given in Eq. (17), the fit parameter $B$ should be compared against $\beta/\nu-d$, to obtain $\beta=0.119(93)$ which compared against its exact value of $0.125$ for the 2D Ising model, gives an estimate for the exponent. Finally, we show the results of the fit of $H_{0}$ we obtained for each lattice size, plotted against $L^{\beta/\nu-d}$ in Fig. 10.
Figure 10: Finite size scaling of $\mbox{Im}(H_{0})$. To guide the eye, we plot data versus $L^{1/8-2}$, where the correct critical exponents $\beta$ and $\nu$ are taken. The value obtained from fits is $\beta/\nu-d=-\beta\delta=-1.881(93)$, as shown in Table 1, which also gives $\beta=0.119(93)$.
In principle one should be able to follow these steps to estimate the critical region for QCD using Taylor expansions from lattice QCD. The relevant parameters which control the critical region in QCD will be the baryon density. We have tried to make some concrete steps in this direction recently in [64]. However, being a much more complicated theory, both numerically and conceptually, we may have to wait for some time to be able to do that.
@@
IIa $\mathbf{0.4404(19)}$ $-0.216(70)$ $\beta_{c}\sim 0.4407$ $1.44$
III $1.30(24)$ $\mathbf{1.881(93)}$ $d-\beta/\nu=\beta\delta=1.875$ $1.2$
Table 1: Results for the fits shown in Eqs. (12,13,16), shown for each row respectively. For Fit I, $B$ has to be compared with the exact value stated, whereas for Fit II, the intercept gives $\beta_{c}$, hence the exact value has to be compared with $A$. For the last fit, Fit III, the fit parameter $B$ has to be compared to the exact value of the critical exponent product, namely, $\beta\delta$.
@@ VI Conclusions and Outlook
As a first step we simulated the 2D Ising model using a cluster spin flip algorithm in two ways. For the LY zero analysis, we simulated the model on varying lattice sizes at a set of values of the external magnetic field. These simulations were performed at $T_{c}$, and the goal was to approximate the average magnetization as a rational function of the external magnetic field and study the structure of zeros and poles that arise. On the one hand we were able to verify numerically, many properties of the LY zeros including the famous circle theorem for the Ising model, observing that only the genuine poles (un-cancelled and stable) of magnetization lie on the purely imaginary $H$ axis. It was further observed that the number of genuine poles poles increases with volume and for simulations at $T_{c}$, comes closer to the real $H$ axis. In order to verify that these were indeed physical effects, volume scaling of the zeros, using the prescription in [11, 22] were performed leading to a decent estimate of the combination of critical exponents $\beta\delta=1.881(93)$. The fact that LY zeros can be studied at $T_{c}$ is in our approach fully self-consistent. In fact, Fisher zeros were also studied by approximating the specific heat with a multi-point Padé function and studying its poles in the complex $\beta$ plane. Finite size scaling of these zeros following the prescription of [22] was done to obtain again, precise values of the critical exponent $\nu=1.014(60)$ and of the critical inverse temperature $\beta_{c}=0.4404(19)$. These results give us some confidence in our pursuits of studying the QCD phase diagram using lattice QCD simulations combined with multi-point Padé method. The main caveat being that in the case of the Ising model it was not very computationally expensive to reach statistics of the order of $\sim 625$K configurations for each lattice, for each value of $H$ and $\beta$. These kind of statistics are not currently realistic for lattice QCD simulations.
Another direction to pursue would be to study the universal location of the Lee-Yang edge singularities as done using the Functional Renormalisation Group (FRG) approach in [65, 66] or using a suitable parametrization of lattice data as done in [67] for the $O(N)$ models. In order to do this using our approach, we would need accurate values of the closest LY zero to the real $H$ axis, but at temperatures $T>T_{c}$. We would like to close by inviting the reader to use the multi-point Padé method for extracting Lee-Yang and Fisher zeros in their choice of models. All data from our calculations, presented in the figures of this paper, can be found in [68].
@@ Acknowledgements
S.S. would like to thank Christian Schmidt, Frithjof Karsch and Vladimir Skokov for interesting discussions on the critical phenomena and Lee-Yang zeros. This work was supported by the (i) European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie Grant Agreement No. H2020-MSCAITN-2018-813942 (EuroPLEx), (ii) Part of this work was done with S.S. being supported by PUNCH4NFDI consortium supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 460248186 (PUNCH4NFDI), (iii) I.N.F.N. under the research project i.e. QCDLAT. A good fraction of this research used computing resources made available through the University of Parma on the UNIPR HPC facility.